OpenStudy (anonymous):

11,42,93,65,57,69,..... what is the 7th number ?explain also.

OpenStudy (anonymous):

where do people get these sequences from?

OpenStudy (anonymous):

i cant say ive seen a sequence on this site that has made any sense in the least bit.

OpenStudy (anonymous):

lol

OpenStudy (akshay_budhkar):

sri is it an olypiad question?

OpenStudy (anonymous):

what is it.

OpenStudy (anonymous):

no.akshay.

OpenStudy (anonymous):

if its a polynomal in n, then its at least a cubic >.>

OpenStudy (anonymous):

I showed last time i can make an infinite number of formulas fitting that pattern using 6 random numbers <.<

OpenStudy (akshay_budhkar):

lol1!!

OpenStudy (anonymous):

last time??

OpenStudy (anonymous):

we get questions like these a lot. someone had posted something similar, and just like this one, we couldnt find a pattern. So using the fact that we never know how to the tail of a sequence acts, i made up a tail, and found a formula. if its correct or not, we will never know...but it does work for the first couple of numbers.

OpenStudy (anonymous):

what is the formula?

OpenStudy (anonymous):

give me 6 numbers between -5 and 5, and not 0.

OpenStudy (anonymous):

its going to look ugly btw, but im 100% sure it works.

OpenStudy (anonymous):

-2,-3,1,2,4,5

OpenStudy (anonymous):

k, give me a couple of mins, gotta do calculations.

OpenStudy (anonymous):

ok, the formula is: \[-\frac{3}{10}(-3)^n+\frac{5}{504}(-2)^n-\frac{335}{9}(1)^n+\frac{1107}{40}(2)^n-\frac{101}{36}(4)^n+\frac{9}{14}(5)^n\] for n = 1,2,3...,6 it equals those numbers up top. http://www.wolframalpha.com/input/?i=%E2%88%923%2F10%28%E2%88%923%29^n%2B5%2F504%28%E2%88%922%29^n%E2%88%92335%2F9%281%29^n%2B1107%2F40%282%29^n%E2%88%92101%2F36%284%29^n%2B9%2F14%285%29^n%2C+n+%3D+6 where you type in the equation, if you scroll to the right you can change the value of n to make sure it works.